OpenStudy (thebadyboy):

Which equations have a positive solution? Choose all answers that are correct.

OpenStudy (anonymous):

For option A, \[-\frac{ w }{ 9 }=8\]\[-w=72\]\[w=?\]

OpenStudy (anonymous):

For option B, \[-12=\frac{ v }{ -3 }\]\[-12\times-3=v\]\[v=?\]

OpenStudy (anonymous):

For option C, \[-4=-\frac{ 2z }{ 5 }\]\[-4\times5=-2z\]\[-20=-2z\]\[\frac{ -20 }{ -2 }=z\]\[z=?\]

OpenStudy (anonymous):

For option D, \[\frac{ x }{ 7 }=-2\]\[x=7\times-2\]\[x=?\]

OpenStudy (anonymous):

@Thebadyboy